# Exploiting Negative Capacitance for Unconventional Coulomb Engineering

### v1

@@ Abstract
It is known that the many-body ground state of a two-dimensional electron system can be tuned through Coulomb engineering by controlling the permittivity of the surrounding media. However, permittivities are traditionally restricted to positive values. In this paper we argue that the negative capacitance effect demonstrated in appropriately engineered structures can open new vistas in Coulomb engineering. Negative permittivities transform the natural repulsive interaction of electrons into an attractive one raising the intriguing possibility of a superconducting ground state. Using models of two-dimensional electron systems with linear and parabolic dispersion relations coupled to environments with negative capacitance, we estimate the strength and sign of the engineered Coulomb interaction and outline parameter regimes that could stabilize correlated electronic phases.
@@ I Introduction
Coulomb interactions between electrons govern a variety of many-body phenomena, from stabilizing interacting phases to shaping the electrical and optical responses of matter. Consequently, several approaches have been developed to highlight the effects driven by Coulomb interactions, bringing them to the forefront of condensed matter physics. For example, one prominent approach [bistritzer2011moire] focuses on revealing Coulomb interactions by suppressing competing kinetic energies in materials engineered to host flat energy bands. This flat-band engineering technique has successfully stabilized interacting phases ranging from superconductivity [cao2018, lu2019superconductors, balents2020superconductivity, oh2021] to the fractional quantum anomalous Hall state [sharpe2019qahe, serlin2020qahe, li2021qahe] in two-dimensional materials [kennes2021moirereview]. However, because this approach does not directly modify Coulomb interactions, the resulting interacting phases are typically stabilized only at low temperatures.
A complementary strategy—known as Coulomb engineering—enables the direct tuning of Coulomb interactions by designing heterostructures with tailored electromagnetic environments [raja2017, song2018]. Although Coulomb engineering has successfully altered the electrical and optical responses of materials [jang2008graph, hwang2012fvr, yu2013qcap, kajino2019coulomb, whelan2020fvr], it remains challenging to tune Coulomb interactions over a wide range and to design interaction-driven phases of matter within this framework. This limitation arises because conventional electromagnetic environments explored so far are largely constrained by the requirement that their static dielectric function (permittivity) must remain positive and be greater than that of vacuum.
@@
This work is motivated by recent experimental demonstrations of Negative Capacitance (NC) [zubko2016, yadav2019, das2021local]. We propose to modify the structures in those works to embed two-dimensional electron systems (2DES) such that the normally repulsive interaction between electrons is rendered attractive, thereby unveiling a new regime of unconventional Coulomb engineering.
The space of applications unlocked by this extended tuning of Coulomb interactions spans both quantitative modifications and qualitative transformations. On the quantitative side, one can expect changes to transport properties and optical response in the normal state of 2DES, potentially exceeding limits of conventional Coulomb engineering [jang2008graph, hwang2012fvr, yu2013qcap, raja2017, song2018, kajino2019coulomb, whelan2020fvr]. On the qualitative side, entirely new phases of matter may emerge [van2023cemott], especially once the customarily repulsive electron–electron interactions are rendered attractive. Of particular interest is superconductivity, which requires the formation of paired electronic states [bcs1957]. We provide estimates of pairing strength through the dimensionless parameter $\lambda-\mu^{*}$, adapted from electron–phonon theory [tolmachev1961, morelanderson, colemanbook], and find that values $\gtrsim 0.1$ are achievable in realistic regimes. Crucially, this parameter is highly tunable within our proposal, suggesting that NC-enabled Coulomb engineering could provide a new route toward stabilizing superconductivity and other ordered electronic phases, potentially even at elevated temperatures.
II Central Idea
@@
To illustrate the main concept, we begin with the idealized structure shown schematically in Fig. 1(a) (connections to real material platforms are made in later sections). A two-dimensional electron system (2DES) is sandwiched between a gated conventional dielectric (DE) of thickness $L_{d}$ and a gated NC material of thickness $L_{nc}$. We model the DE and NC regions with diagonal permittivity tensors $\varepsilon^{nc(d)}=\text{diag}(\varepsilon^{nc(d)}_{\perp},\varepsilon^{nc(d)}_{\perp},\varepsilon^{nc(d)}_{z})$, where $z$ denotes the stacking direction and the superscripts $nc$ and $d$ represent NC and DE regions, respectively.
In the 2DES plane, the static Coulomb interaction $V_{\text{eff}}$ as a function of in-plane wavenumber $q$ is given by
$V_{\text{eff}}(q)=\frac{e^{2}}{2\varepsilon_{0}q}\cdot\frac{1}{\varepsilon_{\text{eff}}(q)}=\frac{v_{q}}{\varepsilon_{\text{eff}}(q)},$ (1)
where the effective dielectric function $\varepsilon_{\text{eff}}(q)$ is obtained self-consistently as
@@
  • The expression for $\varepsilon_{\text{env}}(q)$ reflects the “anisotropic gate screening” effect, derived from Poisson’s equation (see Appendix A).
  • For the electron polarization function $\Pi_{el}$, we use the static limit of the Lindhard function for a 2DES [giuliani2008quantum], which reduces to the Thomas-Fermi (TF) form for wavenumber $q<2k_{F}$—the range relevant for this work.
Figure 1: Central idea. (a) System schematic - a two-dimensional electron system (2DES) surrounded by dielectric (conventional) and negative capacitance (NC) media, with respective thicknesses $L_{d}$ and $L_{nc}$. When the NC material is a ferroelectric, this configuration will be referred to as a Metal-Ferroelectric-2DES-Insulator-Metal (MF2IM) structure. (b) Tunability of the engineered 2DES Coulomb interaction $V_{\text{eff}}$ in the long-wavelength approximation as per Eq. (3), shown as a function of geometric capacitances $C_{nc}$ and $C_{d}$, and quantum capacitance $C_{q}$. The dashed black line indicates the limit of conventional Coulomb engineering ($C_{nc}>0$, region I). Notably, $V_{\text{eff}}$ can become negative (region III) if $C_{nc}$ is sufficiently negative.
Long-wavelength limit: Introducing dimensionless anisotropy factors $\eta_{nc}=\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$ and $\eta_{d}=\sqrt{\varepsilon^{d}_{\perp}/\varepsilon^{d}_{z}}$, we illustrate the central idea of NC-enabled unconventional Coulomb engineering by first considering the limit where $q\eta_{nc}L_{nc}\ll 1$ and $q\eta_{d}L_{d}\ll 1$. In this regime, the wavelengths associated with low-energy scattering of electrons in the 2DES are large compared to the gate distances. The assumption of $q$-independent permittivities in this section additionally requires that this wavelength is large compared to characteristic length scales of the polarizable DE and NC media (elaborated in later sections). Under these conditions, Eq. (2) reduces to a simple form:
@@
where $C_{nc}$ and $C_{d}$ are geometric capacitances per unit area, and $C_{q}$ is the quantum capacitance of the 2DES:
$C_{nc}=\frac{\varepsilon_{0}\varepsilon^{nc}_{z}}{L_{nc}}\hskip 5.69046pt;\hskip 5.69046ptC_{d}=\frac{\varepsilon_{0}\varepsilon^{d}_{z}}{L_{d}}\hskip 5.69046pt;\hskip 5.69046ptC_{q}=2\varepsilon_{0}k_{\text{TF}}.$ (4)
Throughout this work, we set the ‘background’ dielectric constant of 2DES to unity for simplicity.
Eq. (3) encapsulates our central thesis: the inclusion of $C_{nc}<0$ enables the design of structures where $V_{\text{eff}}$ can be negative and engineered to have large magnitude values through careful matching of the constituent capacitances.
Before discussing the scope of this result, we first present a stability condition which constrains the values $C_{nc}$ can assume in the proposed configuration.
Structural Stability: The NC state of a material in isolation is thermodynamically unstable, but as argued in Ref. [salahuddindatta2008], it can be locally stabilized in a constituent layer provided the composite structure is stable against charge fluctuations.
We refer to the structure in Fig. 1(a) as Metal-Ferroelectric-2DES-Insulator-Metal (MF2IM) configuration (specifying NC material as ferroelectric to connect with notation used in literature [park2019modeling, luk2022ferroelectric]). The necessary and sufficient condition for stability of NC in the MF2IM configuration is (assuming $C_{nc}<0$ and $C_{d},C_{q}>0$, see Appendix B for details):
@@
Setting $C_{q}=0$ recovers the familiar stability condition for the widely studied Metal-Ferroelectric-Insulator-Metal (MFIM) structure [salahuddindatta2008, park2019modeling, iniguez2019ferroelectric]. In the opposite limit of $C_{q}\to\infty$, the MF2IM structure decouples into two separate capacitors (assuming the 2DES is connected to a charge reservoir), and NC cannot be stabilized in this case.
Scope: Fig. 1(b) illustrates the scope of Eq. (3) by treating the capacitances as free variables. In conventional Coulomb engineering where all geometric capacitances are positive, the tunability of $V_{\text{eff}}$ is limited to region I (assuming $C_{q}>0$). The inclusion of $C_{nc}<0$ opens up new regimes, indicated as regions II and III. Most strikingly, in region III the sign of $V_{\text{\text{eff}}}$ can be reversed while still satisfying the stability condition in Eq. (5), transforming the naturally repulsive electron-electron interaction into an effective attraction.
Figure 2: Critical length scales. (a) Definition of two regimes based on three length scales. The wavelength of low-energy electron scattering in the 2DES relative to $2d$ (twice the domain width) in ferroelectric periodic domain texture (PDT) governs whether dispersion $\varepsilon^{f}_{z}(q)$ is important in extrinsic NC systems ($2d$ should be replaced by an appropriate lattice-related length scale in the case of intrinsic NC) and relative to $|\eta_{f}|L_{f}$ determines whether the linear approximation of gate screening effect is valid. (b), (c) Illustration of effective Coulomb interaction energy $V_{\text{eff}}$ as a function of wavenumber $q$ as per Eqs. (1) and (2), shown for linear and parabolic 2DES with typical $v_{F}$ and $m^{*}$ values. In both cases, 2DES carrier density $n$ was varied to access the long-wavelength (low $n=10^{15}/\text{m}^{2}$) and short-wavelength (high $n=10^{17}/\text{m}^{2}$) regimes. Note that changing $L_{f}$ (fixed at $4~\text{nm}$ for this figure) is an alternate way to access these regimes.
III Model details
In this section, we provide details of the models we use to estimate parameters appearing in Eq. (2), treating the 2DES carrier density $n$ and the NC material thickness $L_{nc}$ as tunable variables.
NC material: Within the growing literature on negative capacitance, it is useful to distinguish between transient and stabilized NC. Transient NC, such as that reported in [khan2015transient, khan2017transient, hoffmann2018transient, hoffmann2019unveiling, kim2019transient, qiao2021both], occurs during polarization switching and is therefore a non-equilibrium associated with dynamical relaxation of the order parameter. In contrast, stabilized NC [appleby2014stable, gao2014stable, zubko2016, hoffmann2018stabilization, yadav2019, das2021local, qiao2021both, qiao2024stable] refers to the quasi-static, near-equilibrium regime, where the system exhibits static negative permittivity. Beyond this dynamical classification, NC systems can also be categorized by material class— perovskite (such as [yadav2019]), oxide (such as [hoffmann2022intrinsic]), van der Waals (such as [wang2019vdw])—and by the underlying ferroic order, with both ferroelectrics [iniguez2019ferroelectric] and antiferroelectrics [qiao2021both, hoffmann2022antiferroelectric] being reported to exhibit NC behavior.
In the remainder of this work we restrict attention to stabilized NC, although the idea outlined in the previous section could be explored in the transient case as well. Within stabilized NC, two classes of systems are often distinguished: (i) extrinsic NC [zubko2016, yadav2019, das2021local], where the phenomenon is associated with multi-domain configurations and domain wall motion [junquera2023topological], and (ii) intrinsic NC [hoffmann2022intrinsic], where the effect appears to be independent of specific domain configurations. The theory developed in the previous section and summarized in Fig. 1 is formulated without assuming either classification, but we adopt a domain-provided model for extrinsic NC when performing calculations for specific material systems (results in Figs. 2 and 3).
@@ III Model details
We now provide details about the domain-provided model for stabilized NC in perovskite ferroelectric $\text{PbTiO}_{3}$. This choice is motivated by substantial evidence from experiments [zubko2016, yadav2019, das2021local] and corresponding theoretical explanation [luk2018prb, luk2019harnessing] providing a framework for modeling near-equilibrium NC. In ferroelectric $\text{PbTiO}_{3}$ ($\varepsilon^{nc}_{z(\perp)}\to\varepsilon^{f}_{z(\perp)}$, $L_{nc}\to L_{f}$), the polarization response is highly anisotropic, with $\varepsilon^{f}_{z}<0$ and $\varepsilon^{f}_{\perp}>0$. Starting from a periodic domain texture (PDT) as the equilibrium polarization configuration, the static negative permittivity $\varepsilon^{f}_{z}$ is explained as an overscreening effect due to the role of depolarization field in this system [luk2018prb, luk2019harnessing]. Assuming small deviations from equilibrium, the PDT dynamics can be described by an oscillator model for domain wall displacements, where $P$ is the electric polarization and $E_{\text{tot}}$ is the total electric field (external + depolarization):
$\ddot{P}(t)+\Gamma\dot{P}(t)+(\omega_{0}^{2}-\Omega^{2})P(t)=\Omega^{2}\cdot\varepsilon_{0}\varepsilon^{f}_{z,hf}E_{\text{tot}}(t).$ (6)
Here, $\Gamma$ is the damping constant for domain wall motion. The restoring force originates from Coulomb energy associated with excess surface charges when the PDT is displaced from equilibrium. Fourier analysis of the domain structure within Kittel model [kittel1946, junquera2023topological] yields the following expressions for frequency parameters [sidorkin2006domain, luk2018prb]:
Here, $\omega_{0}$ is the characteristic oscillation frequency of the PDT system and $\Omega$ represents the strength of coupling to external electric field.
@@
In a more detailed model, these frequency parameters should exhibit dispersion $\omega_{0}(q),\Omega(q)$ with additional indices for excitation type (longitudinal vs transverse) and branch (acoustic vs optic). The expressions in Eq. (7) account only for the longitudinal optic branch at zero wavenumber (see Appendix C), based on the expectation that this mode contributes dominantly to the NC effect in the long-wavelength limit. We refer to this as the ‘single-mode’ approximation.
Parameters for ferroelectric $\text{PbTiO}_{3}$ are chosen as follows [luk2018prb] : $P_{s}=0.65~\text{Cm}^{-2}$ is the spontaneous polarization, and $M$ is the domain wall mass per unit area estimated using interpolation formula $M=1.3\sqrt{L_{f}[\text{nm}]}\times 10^{-9}~\text{kgm}^{-2}$ [luk2018prb, zhangponomareva]. The equilibrium domain width $d$ is calculated using the Kittel model (explicit formula in Appendix C). $\varepsilon^{f}_{z,hf}=100$ is the “high-frequency” dielectric constant which captures the background polarizability of $\text{PbTiO}_{3}$ arising from other polarization mechanisms in the material, and $\varepsilon^{f}_{\perp}=30$.
@@
$\varepsilon^{f}_{z}\equiv\varepsilon^{f}_{z}(q=0,\omega=0)=\ \varepsilon^{f}_{z,\text{hf}}\cdot\frac{\omega_{0}^{2}}{\omega_{0}^{2}-\Omega^{2}}.$ (8)
Since $\omega_{0}<\Omega$ for ferroelectric $\text{PbTiO}_{3}$ in this model, the above expression yields a negative $\varepsilon^{f}_{z}$.
Dielectric: For the conventional dielectric part of environment (see Fig. 1(a)), we adopt values for hexagonal boron nitride (hBN) as a representative ‘low-k’ dielectric. Specifically, we set $\varepsilon^{d}_{z}=3.4$, $\varepsilon^{d}_{\perp}=6.86$ based on values reported in literature [ohbahbn, laturiahbn, Pierrethbn, kimhbn]. In Fig. 3(b), we choose hafnium oxide ($\text{HfO}_{2}$) as a representative ‘high-k’ dielectric in the long-wavelength regime with $\varepsilon^{d}_{z}=25$. The dielectric thickness is fixed at $L_{d}=4~\text{nm}$ in all cases.
2DES: The bare 2DES (in vacuum, $\varepsilon_{\text{env}}=1$) is modeled as a single-valley non-interacting electron gas, with either linear or parabolic low-energy dispersion. The two cases are characterized respectively by a Fermi velocity $v_{F}$, or an effective mass $m^{*}$. Spin and valley degeneracies are set to $g_{S}=2,g_{V}=1$ respectively. The quantum capacitance $C_{q}$ and Thomas-Fermi screening wavenumber $k_{\text{TF}}$ are
$\begin{aligned} \text{Linear:}\hskip 5.69046ptC_{q}=2\varepsilon_{0}k_{\text{TF}}&=\frac{e^{2}k_{F}}{\pi\hbar v_{F}},\\ \text{Parabolic:}\hskip 5.69046ptC_{q}=2\varepsilon_{0}k_{\text{TF}}&=\frac{e^{2}m^{*}}{\pi\hbar^{2}}.\end{aligned}$ (9)
Note that in the linear case $C_{q}$ and $k_{\text{TF}}$ scale with carrier density as $E_{F}$ ($\propto\sqrt{n}$). In the parabolic case, $C_{q}$ and $k_{TF}$ are independent of $n$.
IV Results
In this section, we apply the models described above to identify critical length and energy scales in the composite structure and estimate the strength of engineered Coulomb interactions in the low-energy regime.
Length scales: Let us first note that the quantity $\eta_{nc}\equiv\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$ is imaginary for an anisotropic NC material like $\text{PbTiO}_{3}$ which has $\varepsilon^{f}_{z}<0~,~\varepsilon^{f}_{\perp}>0$. Consequently the tanhtanh functions in Eq. (2) become tantan functions ($\tanh(ix)=i\tan(x)$):
$\varepsilon_{\text{env}}(q)=\frac{1}{2}\bigg[\frac{|\eta_{f}|\varepsilon^{f}_{z}}{\tan\big(q\hskip 0.85355pt|\eta_{f}|L_{f}\big.)}+\frac{\eta_{d}\varepsilon^{d}_{z}}{\tanh\big(q\hskip 1.42271pt\eta_{d}L_{d}\big.)}\bigg],$ (10)

### v2

@@ Abstract
The many-body ground state of a two-dimensional electron system can be tuned by Coulomb engineering through control of the dielectric environment. However, in conventional dielectrics the static permittivity is restricted to positive values, limiting the accessible interaction regimes. Here we argue that the negative capacitance demonstrated in appropriately engineered structures can open new vistas for Coulomb engineering. The associated negative permittivity could transform the natural repulsive interaction of electrons into an attractive one, raising the intriguing possibility of nontrivial ground states, including superconductivity. Using models of two-dimensional electron systems with linear and parabolic dispersion relations coupled to environments with negative capacitance, we estimate the strength and sign of the engineered Coulomb interaction and outline parameter regimes that could stabilize correlated electronic phases.
@@
Introduction.—Coulomb interactions between electrons govern a variety of many-body phenomena, from stabilizing interacting phases to shaping the electrical and optical responses of matter. Consequently, several approaches have been developed to highlight the effects driven by Coulomb interactions, bringing them to the forefront of condensed matter physics. For example, one prominent approach [5] focuses on revealing Coulomb interactions by suppressing competing kinetic energies in materials engineered to host flat energy bands. This technique has successfully stabilized interacting phases ranging from superconductivity [6, 37, 3, 42] to the fractional quantum anomalous Hall state [54, 53, 35] in two-dimensional materials [28]. A complementary strategy—known as Coulomb engineering—enables the direct tuning of Coulomb interactions by designing heterostructures with tailored electromagnetic environments [49, 56]. Although Coulomb engineering has successfully altered the electrical and optical responses of materials [25, 22, 63, 27, 60], it remains challenging to tune Coulomb interactions over a wide range and to design interaction-driven phases of matter within this framework. This limitation arises because previously explored electromagnetic environments are largely restricted to positive static permittivities. However, on fundamental grounds, environments with local static negative permittivity are permissible [11], raising the question of whether present-day Coulomb-engineered structures can be extended beyond this constraint.
@@
Here we answer this question in the affirmative by proposing and studying novel Coulomb-engineered structures motivated by recent demonstrations of ferroelectric negative capacitance (NC) [65, 61, 10]. Specifically, we propose to modify the structures in those works to embed two-dimensional electron systems (2DES) such that the normally repulsive interaction between electrons is rendered attractive, thereby unveiling a new regime of unconventional Coulomb engineering. On the one hand, this extended tuning of Coulomb interactions can change the transport properties and optical response in the normal state of 2DES, potentially exceeding limits of conventional Coulomb engineering [25, 22, 63, 49, 56, 27, 60]. On the other hand, entirely new phases of matter may emerge [58], especially once the customarily repulsive electron–electron interactions are rendered attractive.
While predicting the ground state is beyond the scope of the present work, a state of particular interest with attractive electron interactions is superconductivity, which requires the formation of paired electronic states [4]. Motivated by this, we estimate the pairing strength in a model platform where attractive interactions are mediated by negative capacitance arising from the motion of ferroelectric domain walls [39]. Within this model, the induced attraction is retarded by the ferroelectric domain wall response, whose characteristic frequencies typically lie well below electronic energy scales, leading to a separation of timescales analogous to the phonon-mediated pairing. We find that the resulting dimensionless pairing parameter $\lambda-\mu^{*}$, adapted from electron–phonon theory [57, 41, 8], can reach values $\gtrsim 0.1$ in realistic experimental regimes. Crucially, by balancing positive and negative capacitances, this parameter is highly tunable within our proposal, suggesting that NC-enabled Coulomb engineering could provide a new route towards designing superconductivity and other ordered electronic phases. More broadly, there has recently been growing interest in cavity-based materials engineering [51, 36], where the collective response of specially designed cavities is leveraged to modify the ground state of solid-state materials. Our results indicate that the collective response of ferroelectrics stabilized in the negative-capacitance regime could offer a useful resource for this paradigm.
@@
Central Idea.—To illustrate the main concept, we begin with the idealized structure shown schematically in Fig. 1(a) (connections to real material platforms are made in later sections). A two-dimensional electron system (2DES) is sandwiched between a gated conventional dielectric (DE) of thickness $L_{d}$ and a gated NC material of thickness $L_{nc}$. We model the DE and NC regions with diagonal permittivity tensors $\varepsilon^{nc(d)}=\text{diag}(\varepsilon^{nc(d)}_{\perp},\varepsilon^{nc(d)}_{\perp},\varepsilon^{nc(d)}_{z})$, where $z$ denotes the stacking direction and the superscripts $nc$ and $d$ represent NC and DE regions, respectively. In the 2DES plane, the static Coulomb interaction $V_{\text{eff}}$ as a function of in-plane wavenumber $q$ is given by:
$V_{\mathrm{eff}}(q)=\frac{v_{q}}{\varepsilon_{\mathrm{eff}}(q)},\qquad v_{q}\equiv\frac{e^{2}}{2\varepsilon_{0}q}.$ (1)
We model the effective dielectric function $\varepsilon_{\mathrm{eff}}(q)$ using linear response polarization functions as: $\varepsilon_{\mathrm{eff}}(q)=1-v_{q}(\Pi_{nc}(q)+\Pi_{d}(q)+\Pi_{el}(q))$. The environment dielectric function $\varepsilon_{\mathrm{env}}(q)\equiv 1-v_{q}(\Pi_{nc}(q)+\Pi_{d}(q))$ can be derived from Poisson’s equation (see Appendix A) and reflects the ‘gate screening’ effect:
@@
Figure 1: Central idea. (a) System schematic: a two-dimensional electron system (2DES) surrounded by dielectric (conventional) and negative-capacitance (NC) media, with respective thicknesses $L_{d}$ and $L_{nc}$. When the NC material is a ferroelectric, this structure is referred to as Metal-Ferroelectric-2DES-Insulator-Metal (MF2IM) configuration. (b) Long-wavelength part of engineered 2DES Coulomb interaction $V_{\text{eff}}$ as per Eq. (3), shown as a function of geometric capacitances $C_{nc}$ and $C_{d}$, and quantum capacitance $C_{q}$. The black dashed line indicates the limit of conventional Coulomb engineering ($C_{nc}>0$, region I). Region II is new but expected to be unstable in the MF2IM configuration. Notably, $V_{\text{eff}}$ can become negative in region III, where thermodynamic stability is satisfied as per Eq. (4).
Introducing dimensionless anisotropy factors $\eta_{nc}=\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$ and $\eta_{d}=\sqrt{\varepsilon^{d}_{\perp}/\varepsilon^{d}_{z}}$, we illustrate the central idea of NC-enabled unconventional Coulomb engineering by first considering the limit where $q\eta_{nc}L_{nc}\ll 1$ and $q\eta_{d}L_{d}\ll 1$. In this regime, the wavelengths associated with low-energy scattering of electrons in the 2DES are large compared to the gate distances and characteristic length scales of the polarizable DE and NC media (elaborated in later sections). For such large wavelengths ($q\to 0$), Eqs. (1)–(2) reduce to a simple form:
@@
where $C_{nc}\equiv\varepsilon_{0}\varepsilon^{nc}_{z}/{L_{nc}}$, $C_{d}\equiv\varepsilon_{0}\varepsilon^{d}_{z}/{L_{d}}$ are geometric capacitances per unit area, and $C_{q}\equiv-e^{2}\Pi_{el}(q\to 0)$ is the quantum capacitance of the 2DES. The expression for $C_{q}$ reflects the compressibility sum rule [16]. For conventional dielectrics, $C_{d}>0$, whereas $C_{nc}$ can be negative for materials operating in the NC regime. Throughout this work we focus on the case of $C_{q}>0$, which may be violated in certain low-density regimes of a 2DES due to exchange and correlation effects [13].
Before discussing the scope of this result, we first present a stability condition which constrains the values $C_{nc}$ can assume in the proposed configuration. The NC state of a material in isolation is thermodynamically unstable, but as argued in Ref. [52], it can be locally stabilized in a constituent layer provided the composite structure is stable against charge fluctuations. We refer to the structure in Fig. 1(a) as Metal-Ferroelectric-2DES-Insulator-Metal (MF2IM) configuration (specifying NC material as ferroelectric to connect with notation used in literature [44, 38]). The necessary and sufficient condition for stability of NC in the MF2IM configuration is (assuming $C_{nc}<0$, and $C_{d},C_{q}>0$, see Appendix B for details):
@@
Setting $C_{q}=0$ recovers the familiar stability condition for the widely studied Metal-Ferroelectric-Insulator-Metal (MFIM) structure [52, 44, 24]. In the opposite limit of $C_{q}\to\infty$, the MF2IM structure decouples into two separate capacitors (assuming the 2DES is connected to a charge reservoir), and NC cannot be stabilized in this case.
Fig. 1(b) illustrates the first main result of this work by treating the capacitances as free variables. In conventional Coulomb engineering, where all capacitances are positive, the tunability of $V_{\text{eff}}$ is limited to region I. The inclusion of $C_{nc}<0$ opens additional regimes, labeled regions II and III. According to Eq. (4), region II is unstable against charge fluctuations. If the system is tuned into this region—for example, by varying $C_{q}$—the charge distribution will reconfigure to restore stability. This behavior may provide a new diagnostic tool for probing NC materials. In contrast, region III remains stable while reversing the sign of $V_{\mathrm{eff}}$, converting the intrinsically repulsive electron–electron interaction into an effective attraction. Moreover, as per Eq. (3), the magnitude of this attractive $V_{\mathrm{eff}}$ can be engineered to reach large values through careful matching of the constituent capacitances. In the remainder of this work, we focus on region III, emphasizing model experimental platforms capable of realizing this regime.
Model Platform.— Here, we focus on ferroelectrics as the specific choice of NC material. Within the growing literature on negative capacitance, one distinguishes transient NC [30, 29, 19, 17, 31, 47], which occurs during polarization switching and reflects nonequilibrium order-parameter relaxation, from stabilized NC [2, 15, 65, 20, 61, 10, 47, 48], corresponding to a quasi-static regime with negative permittivity. NC systems are further classified by material class—perovskites [61], oxides [18], van der Waals materials [59]—and by ferroic order, including ferroelectrics  [24] and antiferroelectrics [47, 21]. Here we focus on stabilized NC, distinguishing extrinsic NC [65, 61, 10], associated with multidomain configurations and domain-wall motion [26], from intrinsic NC [18], which appears independent of domain structure. While the general framework presented in Fig. 1 is agnostic to the underlying mechanism, our material-specific calculations (Figs. 2 and 3) employ a domain-based model of extrinsic NC, which we now detail for stabilized NC in perovskite ferroelectrics, exemplified by $\text{PbTiO}_{3}$. This choice is motivated by substantial evidence from experiments [65, 61, 10] and corresponding theoretical explanation [39, 40] providing a framework for modeling near-equilibrium NC.
@@
In ferroelectric $\text{PbTiO}_{3}$ ($\varepsilon^{nc}_{z(\perp)}\to\varepsilon^{f}_{z(\perp)}$, $L_{nc}\to L_{f}$), the polarization response is highly anisotropic, with $\varepsilon^{f}_{z}<0$ and $\varepsilon^{f}_{\perp}>0$. Starting from a periodic domain texture (PDT) as the equilibrium polarization configuration, the static negative permittivity $\varepsilon^{f}_{z}$ is explained as an overscreening effect due to the role of depolarization field in this system [39, 40]. Assuming small deviations from equilibrium, the PDT dynamics can be described by an oscillator model for domain wall displacements, where $P$ is the electric polarization and $E_{\text{tot}}$ is the total electric field (external + depolarization):
$\ddot{P}(t)+(\omega_{0}^{2}-\Omega^{2})P(t)=\Omega^{2}\varepsilon_{0}\varepsilon^{f}_{z,hf}E_{\text{tot}}(t).$ (5)
@@
The restoring force originates from Coulomb energy associated with excess surface charges when the PDT is displaced from equilibrium. Here, $\omega_{0}$ is the characteristic oscillation frequency of the PDT system and $\Omega$ represents the strength of coupling to external electric field. We set the damping [39] to zero, as we focus on the static response in this work. In a more detailed model, the frequency parameters should exhibit dispersion $\omega_{0}(q),\Omega(q)$ with additional indices for excitation type (longitudinal vs transverse) and branch (acoustic vs optic). The expressions which we use to calculate $\omega_{0},\Omega$ in this work (see Appendix C) account for only the longitudinal optic branch at zero wavenumber, based on the expectation that this mode contributes dominantly to the NC effect in the long-wavelength limit (defined below). We refer to this as the ‘single-mode’ approximation.
Figure 2: Critical length scales. (a) The wavelength of low-energy electron scattering in the 2DES, $(2k_{F})^{-1}$, relative to $2d$ (twice the domain width) in ferroelectric periodic domain texture (PDT) governs whether dispersion $\varepsilon^{f}_{z}(q)$ is important in extrinsic NC case, and relative to $|\eta_{f}|L_{f}$ (scaled ferroelectric thickness) determines whether the linear approximation of gate screening effect is valid. In the intrinsic NC case, $2d$ should be replaced by an appropriate lattice-related length scale. (b), (c) Illustration of effective Coulomb interaction energy $V_{\text{eff}}$ as a function of wavenumber $q$ as per Eqs. (1)–(2), for linear and parabolic 2DES with typical parameter values $v_{F}=1\times 10^{6}~\text{m}/\text{s}$ and $m^{*}=0.1m_{e}$ respectively. Here, 2DES carrier density $n$ was varied to access the long-wavelength (low $n=10^{15}/\text{m}^{2}$) and short-wavelength (high $n=10^{17}/\text{m}^{2}$) regimes for fixed $L_{f}=4~\text{nm}$.
@@
$\varepsilon^{f}_{z}\equiv\varepsilon^{f}_{z}(q=0,\omega=0)=\frac{\varepsilon^{f}_{z,\mathrm{hf}}\,\omega_{0}^{2}}{\omega_{0}^{2}-\Omega^{2}}.$ (6)
Here $\varepsilon^{f}_{z,hf}>0$ denotes the high-frequency dielectric constant (see Appendix C for more details). Since $\omega_{0}<\Omega$ for ferroelectric $\text{PbTiO}_{3}$ in this model, the expression in Eq. (6) yields a negative $\varepsilon^{f}_{z}$.
For the conventional dielectric part of environment (see Fig. 1(a)), we adopt values for hexagonal boron nitride (hBN) as a representative ‘low-k’ dielectric. Specifically, we set $\varepsilon^{d}_{z}=3.4$, $\varepsilon^{d}_{\perp}=6.86$ based on values reported in literature [43, 34, 46, 32]. In Fig. 3(b), we choose hafnium oxide ($\text{HfO}_{2}$) as a representative ‘high-k’ dielectric in the long-wavelength regime with $\varepsilon^{d}_{z}=25$. The dielectric thickness is fixed at $L_{d}=4~\text{nm}$ in all cases.
The bare 2DES (in vacuum, $\varepsilon_{\text{env}}=1$) is modeled as a single-valley non-interacting electron gas, with either linear or parabolic low-energy dispersion. The two cases are characterized respectively by a Fermi velocity $v_{F}$, or an effective mass $m^{*}$. Spin and valley degeneracies are set to $g_{s}=2$, $g_{v}=1$ respectively. For the electronic screening, we use the Thomas-Fermi (TF) model and set $v_{q}\Pi_{el}=-k_{\mathrm{TF}}/q$, where the TF screening wavenumber $k_{\text{TF}}$ [9] is given by:
$k_{\mathrm{TF}}=\frac{e^{2}}{\pi\hbar}\times\begin{cases}k_{F}/v_{F},&\text{Linear},\\ m^{*}/\hbar,&\text{Parabolic}.\end{cases}$ (7)
Neglecting exchange and correlation effects, the compressibility sum rule is satisfied with $C_{q}=2\varepsilon_{0}k_{\mathrm{TF}}$. We set the ‘background’ dielectric constant of 2DES to unity for simplicity.
Results.—We begin by analyzing the Fourier-transformed Coulomb interaction in the model platform introduced above. To this end, we first note that the quantity $\eta_{nc}\equiv\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$ is imaginary for an anisotropic NC material like $\text{PbTiO}_{3}$ which has $\varepsilon^{f}_{z}<0~,~\varepsilon^{f}_{\perp}>0$. Consequently the tanhtanh function in Eq. (2) becomes a tantan function ($\tanh(ix)=i\tan(x)$):
$\varepsilon_{\mathrm{env}}(q)=\frac{1}{2}\bigg[\frac{|\eta_{f}|\varepsilon^{f}_{z}}{\tan\big(q\hskip 0.85355pt|\eta_{f}|L_{f}\big)}+\frac{\eta_{d}\varepsilon^{d}_{z}}{\tanh\big(q\hskip 1.42271pt\eta_{d}L_{d}\big)}\bigg],$ (8)